# Composite Flow Matching for Reinforcement Learning with Shifted-Dynamics Data

### v3

@@ Appendix E Algorithm of CompFlow built on Soft-Actor-Critic
We begin by reparameterizing the expectation using the definition of the pushforward distribution induced by the source flow, $p_{1}=\psi_{\theta}(x,1)\#p_{0}$. Specifically,
$\lim_{k\to\infty}\mathbb{E}_{x_{0}\sim p_{0}}\left[\left\|\psi_{\theta}(x_{0},1)-\psi^{k}_{\phi}\left(\psi_{\theta}(x_{0},1),2\right)\right\|^{2}\right]=\lim_{k\to\infty}\mathbb{E}_{x_{1}\sim p_{1}}\left[\left\|x_{1}-\psi^{k}_{\phi}(x_{1},2)\right\|^{2}\right].$
According to Proposition 3.1 in [generale2024conditional], as $\eta\to\infty$, mass transport in the conditional variables $(s,a)$ vanishes. That is, the transport is concentrated exclusively in the next-state space, and optimal transport is performed independently within each fixed $(s,a)$ pair when $\eta\rightarrow\infty$.
Therefore, for any fixed $(s,a)$, we can invoke Theorem 4.2 from [pooladian2023multisample], which guarantees that the expected transport cost under the batch-trained flow $\psi^{k}_{\phi}$ converges to the squared 2-Wasserstein distance between $p_{1}$ and $p_{2}$ as the batch size $k\to\infty$. Hence,
$\lim_{k\to\infty}\mathbb{E}_{x_{1}\sim p_{1}}\left[\left\|x_{1}-\psi^{k}_{\phi}(x_{1},2)\right\|^{2}\right]=W_{2}^{2}(p_{1},p_{2}),$
which concludes the proof. ∎
@@
We use the MuJoCo datasets from D4RL [fu2020d4rl] as our offline data. These datasets are collected from continuous control environments in Gym [brockman2016openai], simulated using the MuJoCo physics engine [todorov2012mujoco]. We focus on three benchmark tasks: HalfCheetah, Hopper, and Walker2d, and evaluate across three dataset types: medium, medium-replay, and medium-expert.
@@
We use Kinematic Shift Tasks from the benchmark [lyu2024odrl]. We select most shift level ’hard’ to make the tasks more challenging
@@
We use Morphology Shift Tasks from the benchmark [lyu2024odrl]. We select most shift level ’hard’ to make the tasks more challenging
@@
Following [lyu2024odrl], the friction shift is implemented by altering the $\mathsf{friction}$ attribute in the $\mathsf{geom}$ elements. The frictional components are adjusted to 5.0 times the frictional components in the offline environment. The following is an example for the Hopper robot.
@@
where $\lambda=\frac{\omega}{\frac{1}{N}\sum_{(s_{j},a_{j})}\min_{i=1,2}Q_{\varsigma_{i}}(s_{j},a_{j})}$ and $\omega\in\mathbb{R}^{+}$ is a normalization coefficient. We train BC-SAC for 400K gradient steps, collecting online data every 10 steps. We use the hyperparameters recommended in [lyu2024odrl].
H2O: H2O [niu2022trust] trains domain classifiers to estimate dynamics gaps and uses them as importance sampling weights during critic training. It also incorporates a CQL loss to encourage conservatism. Since the original H2O is designed for the Online-Offline setting, we adapt the objective to the Offline-Online setting. The critic loss is:
@@
BC-VGDF: BC-VGDF [xu2023cross] filters offline transitions whose estimated values align closely with those from the online environment. It trains an ensemble of dynamics models to predict next states from raw state-action pairs under the online dynamics. Each predicted next state is evaluated by the policy to obtain a value ensemble $\{Q(s^{\prime}_{i},a^{\prime}_{i})\}_{i=1}^{M}$, forming a Gaussian distribution. A fixed percentage ($\xi\%$) of offline transitions with the highest likelihood under this distribution are retained. The critic loss is:
@@
BC-PAR: BC-PAR [lyu2024cross] addresses dynamics mismatch through representation mismatch, measured as the deviation between the encoded source state-action pair and its next state. It employs a state encoder $f_{\psi}$ and a state-action encoder $g_{\xi}$, both trained on the target domain. The encoder loss is:
@@
CompFlow. When training the target flow, we use a quadratic cost function and employ the Python Optimal Transport (POT) library [flamary2021pot] to compute the optimal transport plan for each minibatch using the exact solver. Additional hyperparameters are provided in Table 3 and Table 4. Since the exploration bonus term is closely tied to properties of the environment—such as the state space, action space, and reward structure—it is expected that the optimal exploration strength $\beta$ varies across tasks. We perform a sweep over $\beta\in\{0.01,0.1,0.2\}$, and select the offline data selection ratio $\xi\%$ from $\{30\%,50\%\}$.
@@
Optimizer Adam [kingma2015adam]
@@
We use the green security simulator in [xu2021robust]. The model is a Markov decision process with state, action, transitions, and a terminal return. We summarize the parts needed to reproduce our experiments.

### v4

@@ I.2 Proof
We prove (47) by matching an upper bound and a lower bound, which also implies that any limiting optimal coupling concentrates on the diagonal set $\{y=y^{\prime}\}$. Finally, we connect the resulting conditional OT structure to OT-FM trained with the augmented matching cost (44).
Step 1 (Upper bound via a diagonal coupling). Since $\alpha$ and $\beta$ share the same $\mathcal{Y}$-marginal $\mu$, we may couple them by matching the conditioning variable $y$ exactly. For $\mu$-a.e. $y$, let $q_{y}^{\star}\in\Pi(\alpha_{y},\beta_{y})$ be an optimal coupling achieving $W_{2}^{2}(\alpha_{y},\beta_{y})$. A measurable selection $y\mapsto q_{y}^{\star}$ can be chosen (see, e.g., [63, Chapter 5] or [52, Section 1.4]). Define $\bar{\gamma}\in\Pi(\alpha,\beta)$ by
$\bar{\gamma}(dy,dx,dy^{\prime},dx^{\prime}):=\mu(dy)\,q_{y}^{\star}(dx,dx^{\prime})\,\delta_{y}(dy^{\prime}),$ (49)
where $\delta_{y}$ is the Dirac measure at $y$. Because $\bar{\gamma}$ is supported on $\{y=y^{\prime}\}$, for every $\eta>0$,
$\int c_{\eta}\,d\bar{\gamma}=\int_{\mathcal{Y}}\int_{\mathcal{X}\times\mathcal{X}}\|x-x^{\prime}\|_{2}^{2}\,dq_{y}^{\star}(x,x^{\prime})\,\mu(dy)=\int_{\mathcal{Y}}W_{2}^{2}(\alpha_{y},\beta_{y})\,\mu(dy).$
Therefore,
$\limsup_{\eta\to\infty}W_{2,\eta}^{2}(\alpha,\beta)\leq\int_{\mathcal{Y}}W_{2}^{2}(\alpha_{y},\beta_{y})\,\mu(dy).$ (50)
Step 2 ($y$-transport vanishes as $\eta\to\infty$). Let $\gamma_{\eta}$ be an optimal coupling attaining $W_{2,\eta}^{2}(\alpha,\beta)$ and define
$C_{\star}:=\int_{\mathcal{Y}}W_{2}^{2}(\alpha_{y},\beta_{y})\,\mu(dy)<\infty.$
By Step 1, $W_{2,\eta}^{2}(\alpha,\beta)\leq C_{\star}$, hence
$\int c_{\eta}\,d\gamma_{\eta}=W_{2,\eta}^{2}(\alpha,\beta)\leq C_{\star}.$
Since $c_{\eta}((y,x),(y^{\prime},x^{\prime}))\geq\eta\|y-y^{\prime}\|_{2}^{2}$, it follows that
$\int\|y-y^{\prime}\|_{2}^{2}\,d\gamma_{\eta}\leq\frac{C_{\star}}{\eta}.$ (51)
Step 3 (Extract a weak limit supported on the diagonal). Because $\alpha$ and $\beta$ have finite second moments, $\Pi(\alpha,\beta)$ is tight, hence $\{\gamma_{\eta}\}_{\eta>0}$ is tight. Thus there exist $\eta_{n}\to\infty$ and $\gamma_{\infty}$ such that $\gamma_{\eta_{n}}\Rightarrow\gamma_{\infty}$ weakly. Let $f(y,y^{\prime}):=\|y-y^{\prime}\|_{2}^{2}$ and define the bounded continuous truncation $f_{M}:=\min\{f,M\}$. Since $f_{M}$ is bounded and continuous,
$\int f_{M}\,d\gamma_{\eta_{n}}\to\int f_{M}\,d\gamma_{\infty}.$
Moreover, $0\leq f_{M}\leq f$, so by (51),
$\int f_{M}\,d\gamma_{\infty}=\lim_{n\to\infty}\int f_{M}\,d\gamma_{\eta_{n}}\leq\lim_{n\to\infty}\int f\,d\gamma_{\eta_{n}}=0.$
Letting $M\to\infty$ and applying monotone convergence yields $\int f\,d\gamma_{\infty}=0$, which implies $y=y^{\prime}$ holds $\gamma_{\infty}$-a.s. Hence $\gamma_{\infty}$ is supported on $\{(y,x;y,x^{\prime}):\ y=y^{\prime}\}$.
Step 4 (Lower bound and convergence of values). Since $\gamma_{\infty}$ is supported on $\{y=y^{\prime}\}$ and has marginals $\alpha$ and $\beta$, it disintegrates as
$\gamma_{\infty}(dy,dx,dx^{\prime})=\mu(dy)\,\gamma_{y}(dx,dx^{\prime}),$
where $\gamma_{y}\in\Pi(\alpha_{y},\beta_{y})$ for $\mu$-a.e. $y$. Let $g(x,x^{\prime}):=\|x-x^{\prime}\|_{2}^{2}$ and $g_{M}:=\min\{g,M\}$. Because $0\leq g_{M}\leq c_{\eta_{n}}$,
$W_{2,\eta_{n}}^{2}(\alpha,\beta)=\int c_{\eta_{n}}\,d\gamma_{\eta_{n}}\geq\int g_{M}\,d\gamma_{\eta_{n}}.$
By weak convergence, $\int g_{M}\,d\gamma_{\eta_{n}}\to\int g_{M}\,d\gamma_{\infty}$. Letting $M\to\infty$ and using monotone convergence yields
$\liminf_{n\to\infty}W_{2,\eta_{n}}^{2}(\alpha,\beta)\geq\int g\,d\gamma_{\infty}=\int_{\mathcal{Y}}\left(\int\|x-x^{\prime}\|_{2}^{2}\,d\gamma_{y}(x,x^{\prime})\right)\mu(dy)\geq\int_{\mathcal{Y}}W_{2}^{2}(\alpha_{y},\beta_{y})\,\mu(dy).$
Combining this with (50) proves (47). Moreover, the last inequality can only be tight if
$\int\|x-x^{\prime}\|_{2}^{2}\,d\gamma_{y}(x,x^{\prime})=W_{2}^{2}(\alpha_{y},\beta_{y})\quad\text{for $\mu$-a.e.\ $y$},$
which implies $\gamma_{y}$ is a $W_{2}$-optimal coupling between $\alpha_{y}$ and $\beta_{y}$ for $\mu$-a.e. $y$. This proves Claim 2.
Step 5 (OT-FM with augmented matching cost; then $\eta\to\infty$). Fix any conditioning value $y=(s,a)$ and recall that the offline flow satisfies
$\psi^{\rm off}_{\theta}(\cdot,1\mid y)\#p_{0}=\alpha_{y}.$
Let $x_{1}:=\psi^{\rm off}_{\theta}(x_{0},1\mid y)$ with $x_{0}\sim p_{0}$, so that $x_{1}\sim\alpha_{y}$.
(a) OT-FM population limit for a fixed $\eta$. Fix $\eta>0$. Our OT-FM procedure builds minibatch OT matchings between joint samples $(y_{i},x_{i})\sim\alpha$ and $(y^{\prime}_{j},x^{\prime}_{j})\sim\beta$ using the augmented cost $c_{\eta}$, and then trains a conditional map $x\mapsto\psi^{{\rm on},k,\eta}_{\phi}(x,2\mid y)$ by regressing onto the matched $x^{\prime}$ values.
Let $\gamma_{\eta}\in\Pi(\alpha,\beta)$ be an optimal coupling attaining $W_{2,\eta}^{2}(\alpha,\beta)$. Disintegrate $\gamma_{\eta}$ with respect to its first marginal $\alpha$:
$\gamma_{\eta}(dy,dx,dy^{\prime},dx^{\prime})=\alpha(dy,dx)\,\Gamma^{\eta}_{y,x}(dy^{\prime},dx^{\prime}),$
where $\Gamma^{\eta}_{y,x}$ is a probability kernel on $\mathcal{Y}\times\mathcal{X}$. Define the induced conditional law of $x^{\prime}$ given $(y,x)$ by marginalizing out $y^{\prime}$:
$\Lambda^{\eta}_{y,x}(dx^{\prime}):=\int_{\mathcal{Y}}\Gamma^{\eta}_{y,x}(dy^{\prime},dx^{\prime}).$
The associated barycentric projection is
$b_{\eta}(y,x):=\int_{\mathcal{X}}x^{\prime}\,\Lambda^{\eta}_{y,x}(dx^{\prime}).$
Under assumptions (A1)–(A3) and OT-FM convergence guarantees (e.g., Theorem 4.2 of [47], applied with the quadratic cost in the augmented space), training OT-FM to convergence with minibatch size $k$ yields the population limit
$\lim_{k\to\infty}\mathbb{E}_{(y,x)\sim\alpha}\Big[\big\|\psi^{{\rm on},k,\eta}_{\phi}(x,2\mid y)-b_{\eta}(y,x)\big\|_{2}^{2}\Big]=0.$ (52)
In particular, by conditioning on $y$, for $\mu$-a.e. fixed $y$,
$\lim_{k\to\infty}\mathbb{E}_{x\sim\alpha_{y}}\Big[\big\|\psi^{{\rm on},k,\eta}_{\phi}(x,2\mid y)-b_{\eta}(y,x)\big\|_{2}^{2}\Big]=0.$ (53)
(b) Send $\eta\to\infty$ (reduction to conditional OT). Take a subsequence $\eta_{n}\to\infty$ such that $\gamma_{\eta_{n}}\Rightarrow\gamma_{\infty}$ as in Step 3. Let $b_{\eta_{n}}$ be the barycentric projection induced by $\gamma_{\eta_{n}}$ as above. Since $\beta$ has finite second moment, $\{b_{\eta_{n}}\}_{n}$ is uniformly bounded in $L^{2}(\alpha)$, so by extracting a further subsequence (not relabeled) we may assume
$b_{\eta_{n}}\rightharpoonup b_{\infty}\quad\text{weakly in }L^{2}(\alpha).$ (54)
Because $\gamma_{\infty}$ is supported on $\{y=y^{\prime}\}$ and disintegrates as
$\gamma_{\infty}(dy,dx,dx^{\prime})=\mu(dy)\,\gamma_{y}(dx,dx^{\prime}),\qquad\gamma_{y}\in\Pi(\alpha_{y},\beta_{y})\ \text{$W_{2}$-optimal},$
the conditional law of $x^{\prime}$ given $(y,x)$ under $\gamma_{\infty}$ is supported on the conditional coupling $\gamma_{y}$.
Under assumption (A2), for $\mu$-a.e. $y$, the optimal quadratic-cost coupling $\gamma_{y}$ is induced by a Monge map $T_{y}:\mathcal{X}\to\mathcal{X}$ (i.e., $x^{\prime}=T_{y}(x)$ $\gamma_{y}$-a.s.). Hence the limiting barycentric projection satisfies
$b_{\infty}(y,x)=T_{y}(x)\quad\text{for $\alpha$-a.e.\ $(y,x)$}.$
Therefore, for $\mu$-a.e. $y$,
$\mathbb{E}_{x\sim\alpha_{y}}\!\big[\|x-b_{\infty}(y,x)\|_{2}^{2}\big]=\mathbb{E}_{x\sim\alpha_{y}}\!\big[\|x-T_{y}(x)\|_{2}^{2}\big]=W_{2}^{2}(\alpha_{y},\beta_{y}).$ (55)
Finally, combining (53) with the identification $b_{\infty}(y,x)=T_{y}(x)$ and (55), we obtain for $\mu$-a.e. $y$,
$\lim_{\eta\to\infty}\;\lim_{k\to\infty}\;\mathbb{E}_{x\sim\alpha_{y}}\Big[\big\|x-\psi^{{\rm on},k,\eta}_{\phi}(x,2\mid y)\big\|_{2}^{2}\Big]=W_{2}^{2}(\alpha_{y},\beta_{y}).$
Substituting $x=\psi^{\rm off}_{\theta}(x_{0},1\mid y)$ with $x_{0}\sim p_{0}$ yields (48). This completes the proof. ∎
Intuition: Why do we need matched $(s,a)$ marginals? When $(s,a)$ is continuous, it is unlikely to observe many online transitions $s^{\prime}_{\rm on}$ with exactly the same conditioning pair $(s,a)$. As a result, the OT-FM objective must couple two sets of transitions whose conditioning variables differ across samples. To recover the conditional Wasserstein distance $W_{2}\!\left(p_{\rm off}(\cdot\mid s,a),\,p_{\rm on}(\cdot\mid s,a)\right)$ for a fixed $(s,a)$, we add $(s,a)$ into the OT cost: $\|s^{\prime}_{\rm off}-s^{\prime}_{\rm on}\|_{2}^{2}+\eta\!\left(\|s_{\rm off}-s_{\rm on}\|_{2}^{2}+\|a_{\rm off}-a_{\rm on}\|_{2}^{2}\right).$ As $\eta$ grows, transporting mass across different $(s,a)$ becomes prohibitively expensive, so the optimal coupling is forced to match samples with similar $(s,a)$ and transport occurs primarily in the next-state space. This reduction requires the two empirical measures to share the same marginal distribution over $(s,a)$; otherwise, the coupling is forced to move mass in $(s,a)$ even when $\eta$ is large. In practice, we ensure this by sampling $(s_{\rm off},a_{\rm off})$ from the online dataset $\mathcal{D}_{\rm on}$ to construct the offline-generated batch (then sampling $s^{\prime}_{\rm off}\sim p_{\rm off}(\cdot\mid s,a)$ using the offline flow), while sampling online transitions $(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on})$ independently from $\mathcal{D}_{\rm on}$.
@@
We use the MuJoCo datasets from D4RL [16] as our offline data. These datasets are collected from continuous control environments in Gym [4], simulated using the MuJoCo physics engine [62]. We focus on three benchmark tasks: HalfCheetah, Hopper, and Walker2d, and evaluate across three dataset types: medium, medium-replay, and medium-expert.
@@
We use Kinematic Shift Tasks from the benchmark [38]. We select most shift level ’hard’ to make the tasks more challenging
@@
We use Morphology Shift Tasks from the benchmark [38]. We select most shift level ’hard’ to make the tasks more challenging
@@
Following [38], the friction shift is implemented by altering the $\mathsf{friction}$ attribute in the $\mathsf{geom}$ elements. The frictional components are adjusted to 5.0 times the frictional components in the offline environment. The following is an example for the Hopper robot.
@@
where $\lambda=\frac{\omega}{\frac{1}{N}\sum_{(s_{j},a_{j})}\min_{i=1,2}Q_{\varsigma_{i}}(s_{j},a_{j})}$ and $\omega\in\mathbb{R}^{+}$ is a normalization coefficient. We train BC-SAC for 400K gradient steps, collecting online data every 10 steps. We use the hyperparameters recommended in [38].
H2O: H2O [46] trains domain classifiers to estimate dynamics gaps and uses them as importance sampling weights during critic training. It also incorporates a CQL loss to encourage conservatism. Since the original H2O is designed for the Online-Offline setting, we adapt the objective to the Offline-Online setting. The critic loss is:
@@
BC-VGDF: BC-VGDF [65] filters offline transitions whose estimated values align closely with those from the online environment. It trains an ensemble of dynamics models to predict next states from raw state-action pairs under the online dynamics. Each predicted next state is evaluated by the policy to obtain a value ensemble $\{Q(s^{\prime}_{i},a^{\prime}_{i})\}_{i=1}^{M}$, forming a Gaussian distribution. A fixed percentage ($\xi\%$) of offline transitions with the highest likelihood under this distribution are retained. The critic loss is:
@@
BC-PAR: BC-PAR [37] addresses dynamics mismatch through representation mismatch, measured as the deviation between the encoded source state-action pair and its next state. It employs a state encoder $f_{\psi}$ and a state-action encoder $g_{\xi}$, both trained on the target domain. The encoder loss is:
@@
CompFlow. When training the target flow, we use a quadratic cost function and employ the Python Optimal Transport (POT) library [15] to compute the optimal transport plan for each minibatch using the exact solver. Additional hyperparameters are provided in Table 3 and Table 4. Since the exploration bonus term is closely tied to properties of the environment—such as the state space, action space, and reward structure—it is expected that the optimal exploration strength $\beta$ varies across tasks. We perform a sweep over $\beta\in\{0.01,0.1,0.2\}$, and select the offline data selection ratio $\xi\%$ from $\{30\%,50\%\}$.
@@
Optimizer Adam [25]
@@
Figure 7: Well-hidden snares and rangers conducting a patrol to locate them. Photos credit: Uganda Wildlife Authority
We use the green security simulator in [67]. The model is a Markov decision process with state, action, transitions, and a terminal return. We summarize the parts needed to reproduce our experiments.